If 40% of x is equal to 30% of y, what % is x of y?
A. 50%
B. 75%
C. 175%
D. 20%
E. 40%
Hello and please I'm not looking for an answer just a Helpful explanation how can I SOLVE the problem
Thank you,

Answer by richwmiller(17219)

.4x=.3y
.4x/y=.3
x/y=.3./.4
x/y=.75
.75=75%